OpenStudy (anonymous):

Find the limit as x approaches 0 for (x^2-3sinx)/x

OpenStudy (anonymous):

the limit =-3

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ 2x -3cosx}{ 1}\]

OpenStudy (zehanz):

@mathsmind has done it with l'Hôpital's Rule. That's ok. This one could also be done in a more basic way. Split up: \[\lim_{x \rightarrow 0}\frac{ x^2 }{ x }-3\lim_{x \rightarrow 0}\frac{ \sin x }{ x }\] The left one is just\[\lim_{x \rightarrow 0}\frac{ x \cdot x }{ x }=\lim_{x \rightarrow 0}x=0\]The right one is 3 times the well-known standard limit (with an outcome of 1) So the result would be 0-3*1=-3.

OpenStudy (anonymous):

these problem can be worked out just by looking at them

OpenStudy (anonymous):

but it seems even when you tell the rules they repeat same question over and over again

OpenStudy (zehanz):

Just keep explaining will help. Sooner or later, the understanding will arrive...